Compute fxy and fyx for the function f(x,y) = e^(x+y) sinx + 9x^2 + 2xy at (1,2).
1
Expert's answer
2020-04-15T17:24:07-0400

"f_x'=e^{x+y}\sin x+e^{x+y}\cos x+18x+2y" , "f_{xy}''=e^{x+y}\sin x+e^{x+y}\cos x+2"

"f_y'=e^{x+y}\sin x+2x" , "f_{yx}''=""e^{x+y}\sin x+e^{x+y}\cos x+2" then

"f_{xy}''=f_{yx}''=e^{x+y}(\sin x+\cos x)+2\approx29.8" at "(x,y)=(1,2)"